A tube of length L is filled compeletly with an incompressible liquid of mass M and closed at the both ends  the tube is then rotated in horizontal plane about one of its end with a uniform angular velocity W The force exerted by the liquid at the other end is

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

consider a small length dx at adistance x on the rod from the axis of rotation

mass of liquid in dx legth = dm = M/L dx

centrifugal force is acting  on dm mass

 dF = dm x ω2

 F = ∫df  =∫dm x ω2

             =M/L ω2∫x dx        limit 0 to L

             =MLω2/2

this force is balance by normal reaction on fluid by other end

N =   MLω2/2

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Praval Kumar sinha
5 years ago
Look as all the liquid is not concentrated at one place so we can't simply write the formula for centrifugal force . Since we know that the liquid is incompressible .. so it exerts some less force than the centrifugal force acting on mass M .. and hence 
 

consider a small length dx at adistance x on the rod from the axis of rotation

mass of liquid in dx legth = dm = M/L dx

centrifugal force is acting  on dm mass

 dF = dm x ω2

 F = ∫df  =∫dm x ω2

             =M/L ω2∫x dx        limit 0 to L

             =MLω2/2

this force is balance by normal reaction on fluid by other end

N =   MLω2/2

 
hence